OpenStudy (kewlgeek555):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (20 points) @ranga @mathstudent55 @agent0smith @campbell_st @savannaxx_ @DebbieG

OpenStudy (kewlgeek555):

A little bit confused on this one... :-/

OpenStudy (agent0smith):

What :/ this isn't a question.

OpenStudy (kewlgeek555):

It is a question. It is Part A, though...I thought that if guys could help me with Part A, I could do the other ones...

OpenStudy (tkhunny):

Let's see... \(\dfrac{$465}{5}\cdot 7 =\;??\) - -It might be a deal to rent for the whole week!

OpenStudy (kewlgeek555):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points)

OpenStudy (kewlgeek555):

THAT is the whole question @tkhunny and @agent0smith

OpenStudy (agent0smith):

lol that is a question. the first post was not. Always post the whole question.

OpenStudy (kewlgeek555):

Sorry 'bout that. Didn't realize it.

OpenStudy (tkhunny):

I think it was my point that the data are inconsistent. 465/5 = 93 625/7 = 89.2857 So, what's the plan? Does it get cheaper with every day? If so, it will eventually be free!!

OpenStudy (agent0smith):

I'm guessing maybe this is supposed to be a piecewise function?

OpenStudy (kewlgeek555):

Well let me post the whole 3-Part Question. Maybe that will help: Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (agent0smith):

It kinda sounds like a piecewise function... but wtf happens if he rents for 6 days?? 465x if x=<5 625(x/7) if x = 7 465x + 625 if 7 < x =< 12 ... it seems like it's going to continue

OpenStudy (kewlgeek555):

So...it is a sequence??

OpenStudy (agent0smith):

Kinda... but oh i guess at 6 days it's just the daily rate tkhunny worked out... i shouldve used that actually, 93 per day 93x if x=<6 625(x/7) if x = 7 93x + 625 if 7 < x < 14 625(x/7) if x= 14 I guess you can simplify it a bit, no real need for the x/7 business. 93x if x=<6 625 if x = 7 93x + 625 if 7 < x < 14 1250 if x= 14 93x + 1250 if 15 < x < 21 1875 if x= 21 i'm sure there's another way to write this...

OpenStudy (kewlgeek555):

That makes some sense...

OpenStudy (kewlgeek555):

[Sigh]

OpenStudy (kewlgeek555):

How would you put the equation on the graph?

OpenStudy (agent0smith):

Just draw all the segments one at a time, for each of those domains.

OpenStudy (anonymous):

Sorry for the late reply. But I found out the answer. I believe there is a set price before renting. y = 65 + 80x I have no idea how I got it, I was just playing around with the numbers and it popped into my head. I know, weird. But, when we plug in, 5 = x. y = 65 + 80(5) y = 65 + 400 y = 465 When he rents for 5 days, he pays $465. Or when we plug in 7 = x y = 65 + 80(7) y = 65 + 560 y = 625 When he rents for 7 days, he paws $625. I hope that makes sense. ^^

OpenStudy (anonymous):

I need help with these problems too, I still don't understand even though y'all tried to explain them.. Lol.

OpenStudy (tkhunny):

Post on a new thread. We're still here when we can be.

OpenStudy (anonymous):

try dividing the total price by 7?

OpenStudy (anonymous):

@hannanc try dividing the total price by 7